# Rapid and Safe Trajectory Planning over Diverse Scenes through Diffusion Composition

### v1

@@ IV-B3 On the effectiveness of Safety Filter
Compared to RSTP, the F.Rate and C.Rate of RSTP-NSF increase to 4.50% and 11.87%, respectively, emphasizing the necessity of utilizing the safety filter. Additionally, the mean C.T for RSTP-NSF is 0.195 s, comparable to the 0.21 s measured with the safety filter enabled. This indicates minimal computational consumption introduced, thus validating its suitability for time efficiency. Furthermore, the standard deviation of M.TE increases from 1.647 in RSTP to 4.299 in RSTP-NSF, indicating a decrease in weak control feasibility without the safety filter.
(a) Without GSP
(b) With GSP
@@
Figure 4: Diffusion Model Can Ensure Kinematic Constraints. The RSTP method can plan safe trajectories involving large curvature 4a and gear shifting points (GSP) 4b.
Figure 5: Decision-making during Generalization. Illustration of F1TENTH’s velocity of different diffusion models in $CS_{1}$.
IV-C Evaluate Compositional Performance in Real-world
@@
The effectiveness of the model composition for safety is first validated through trajectory planning in real-world scenarios $CS_{1}$ and $CS_{2}$. In scenario $CS_{1}$, the velocities corresponding to trajectories planned by different models are depicted in Fig. 5. The trajectory generated by the DyM1 alone results in a collision with a static obstacle. However, when composing the StM with the DyM1, the StM steers the planned trajectory of DyM1, effectively avoiding collisions with static obstacles. This generalization capability is clearly evident from the decision to decelerate, as marked in Fig. 5. Similarly, in scenario $CS_{2}$ as shown in Fig. 6, the composed model successfully generates a collision-free trajectory. These experimental results demonstrate that the model composition enhances the safety of the RSTP method for real-world applications.
@@
To further evaluate the safety of the proposed RSTP method, three distinct starting points are selected within obstacle-free regions of the map in both scenes $CS_{1}$ and $CS_{2}$. The safety metric used is the minimum distance between the vehicle vertices and both static and dynamic obstacles, with results summarized in Table II. Even with localization uncertainty, the minimum obstacle distance recorded is 0.032 m, confirming adequate safety margins. Moreover, fewer than 10% of trajectory points fall within the defined danger threshold, indicating that only a minor segment of the planned trajectory poses a potential collision risk. Thus, the composed model proves to be sufficiently safe in the context of real-world perception uncertainty.
@@
Figure 6: Safe Planning with Composition. Schematic diagram of trajectories planned by different models in $CS_{2}$.
TABLE II: Real-world Evaluation of Safety Performance of Planned Trajectories Generated by Models Composition.
Scenes Start points D.Thres1 Max Min Mean Std $CS_{1}$ [0.86, 0.03] 6.60% 0.591 0.032 0.283 0.020 [3.03, -0.07] 8.54% 0.559 0.038 0.260 0.020 [2.73, 0.32] 3.77% 0.532 0.054 0.288 0.016 $CS_{2}$ [0.82, 0.07] 4.55% 0.590 0.051 0.278 0.015 [2.76, 0.47] 0.00% 0.593 0.086 0.304 0.019 [2.79, -0.11] 7.41% 0.550 0.036 0.275 0.018
  • 1 The proportion of trajectory points within the Danger Threshold (D.Thres) is calculated as ${N_{\text{Danger}}}/{N_{p}}$, where $N_{\text{Danger}}$ denotes the number of trajectory points whose minimum distances to obstacles fall below the predefined safe threshold (0.08m), and $N_{p}$ represents the number of points.
V Conclusion
@@
This paper introduces the RSTP framework, which combines a diffusion model-based planner with a posterior safety filter. The key contributions include a real-time, fast diffusion planner integrated with onboard perception; a model composition strategy that enables safe generalization to unseen scenarios with high stability; and extensive validation on a real-world F1TENTH vehicle. Experimental results demonstrate the effectiveness of RSTP and highlight the impact of model composition weights on performance. Future work will extend RSTP to racing overtaking scenarios and incorporate constraints-in-the-loop to enhance safety further.
@@
This appendix is structured as follows: Section -A details the experimental setup and hardware configurations; Section -B examines the control feasibility of trajectories generated by the RSTP method; and Section -C evaluates the performance of the model composition within simulation.
-A Experiment Setup
@@
To evaluate the RSTP, training and evaluation datasets are constructed from a real-world $6\,\text{m}\times 6\,\text{m}$ map. From the map’s free navigable area, a comprehensive evaluation dataset $D_{A}$ containing $N_{D}^{\text{all}}=1045$ uniformly sampled initial poses is created. To specifically assess the compositional model, a subset $D_{C}$ with $N_{D}^{\text{comp}}=79$ initial poses is selected from $D_{A}$. In the individual evaluation of the diffusion model, target poses remain fixed. The trajectory horizon $L$ is set to 128 with 100 diffusion steps during training. Denoising employs the DDIM method with 8 steps, and the number of candidate trajectories $N_{\text{filter}}$ for the safety filter is set to 8. Model training and inference are performed using an RTX 4060 GPU and a 4.9 GHz Intel(R) i7-12700 CPU with 48GB RAM.
-B On the Control Feasibility of our RSTP method
@@
The proposed RSTP method also demonstrates excellent control feasibility, as presented in Table I. Specifically, the M.TE of RSTP shows a substantial reduction of 47.23% compared to BLM. This significant improvement indicates that the diffusion model can effectively and implicitly learn vehicle kinematics from the dataset. In contrast, methods such as $\text{A}^{\star}$, $\text{RRT}^{\star}$, and $\text{FMT}^{\star}$, which neglect vehicle kinematic constraints, lead to considerably larger tracking errors. Optimization-based methods, such as RITP and OBCA, achieve notable reductions in M.TE of 70.67% and 79.27%, respectively, yielding trajectories with strong control feasibility. However, this improved accuracy comes at a high computational cost. In comparison, RSTP generates trajectories with strong control feasibility while demands lower computational resources.
-C Effectiveness of Models Composition and Sensitivity Analysis of Compositional Weights
@@
The effectiveness of diffusion models composition for generalization is validated in scenes $CS_{1}$ (Fig. 1) and $CS_{2}$ (Fig. 6), employing four distinct compositional weight configurations, as summarized in Table III. As indicated in the table, the composed model can safely generalize to unseen scenes when appropriate compositional weights are selected, achieving an F.Rate of 0%. However, increasing the weight assigned to conditional models results in a rise in F.Rate across both test scenes. These results demonstrate that the generalization capability is sensitive to compositional weights, and maintaining appropriate weights is critical for minimizing failures.

### v2

@@ V-B3 On the Effectiveness of Safety Filter
Compared to RSTP, the F.Rate and C.Rate of RSTP-NSF increase to 4.50% and 11.87%, respectively, emphasizing the necessity of utilizing the safety filter. Additionally, the mean C.T for RSTP-NSF is 0.195 s, comparable to the 0.21 s measured with the safety filter enabled. This indicates minimal computational overhead introduced, thus validating its suitability for time efficiency. Furthermore, the standard deviation of M.TE increases from 1.647 in RSTP to 4.299 in RSTP-NSF, indicating a decrease in weak kinematic feasibility.
((a))
((b))
@@
Figure 4: Diffusion Model Can Ensure Kinematic Constraints. The RSTP method can plan safe trajectories involving large curvature 4(a) and gear shifting points (GSP) 4(b).
Figure 5: The Composed Model Flexibly Avoids Obstacles in Unseen Scenarios by Adjusting Velocity. Comparison of F1TENTH’s velocity under different diffusion models in $CS_{1}$. The composed model is able to decelerate to avoid static obstacles, whereas the dynamic model (DyM1) fails to avoid the additional static obstacles.
V-C Evaluate Compositional Performance in Real-world
@@
The effectiveness of the model composition for safety is first validated through trajectory planning in the real world. Fig. 5 clearly shows the difference in velocities between the composed model and the individual models. The trajectory generated by the DyM1 alone results in a collision with a static obstacle. However, when composing the StM with the DyM1, the StM steers the planned trajectory of DyM1, effectively avoiding collisions with static obstacles. This generalization capability is clearly evident from the decision to decelerate, as marked in Fig. 5. Fig. 6 presents the collision-free trajectories generated by the composed models in scenes $CS_{1}$ and $CS_{2}$. In both scenarios, the composed models successfully produce trajectories that avoid collisions. These experimental results demonstrate that the model composition enhances the safety of the RSTP method for real-world applications.
@@
To further evaluate the safety of the proposed RSTP method, three distinct starting points are selected within obstacle-free regions of the map in both scenes $CS_{1}$ and $CS_{2}$. The safety metric used is the minimum distance between the vehicle vertices and both static and dynamic obstacles, with results summarized in Table II. Even with localization uncertainty, the minimum obstacle distance recorded is 0.032 m, confirming adequate safety margins. Moreover, fewer than 10% of trajectory points fall within the defined danger threshold, indicating that only a minor segment of the planned trajectory poses a potential collision risk. Thus, the composed model proves to be sufficiently safe in the context of real-world uncertainty.
((a))
((b))
@@
Figure 6: Safe Planning with Composition. Schematic diagram of trajectories planned by different models in composed scenes. Green denotes the ego trajectory; red denotes the dynamic-obstacle trajectory.
TABLE II: The Vehicle–Obstacle Distances in Real-world Execution when Following the Trajectory Generated by the Composed Model.
Scenes Start points Danger1 Max Min Mean Std $CS_{1}$ [0.86, 0.03] 6.60% 0.591 0.032 0.283 0.020 [3.03, -0.07] 8.54% 0.559 0.038 0.260 0.020 [2.73, 0.32] 3.77% 0.532 0.054 0.288 0.016 $CS_{2}$ [0.82, 0.07] 4.55% 0.590 0.051 0.278 0.015 [2.76, 0.47] 0.00% 0.593 0.086 0.304 0.019 [2.79, -0.11] 7.41% 0.550 0.036 0.275 0.018
  • 1 The Danger metric is defined as ${N_{\text{Danger}}}/{N_{all}}$, where $N_{\text{Danger}}$ is the number of points with a minimum distance to other obstacles below 0.08 m, and $N_{all}$ is the total number of points.
VI Conclusion
@@
This paper introduces the RSTP framework, which combines a diffusion model-based planner with a test-time safety filter. The key contributions include a real-time, fast diffusion planner integrated with onboard perception; the MPC-based dataset generation; a model composition strategy that enables safe generalization to unseen scenarios with high stability; and extensive validation on a real-world F1TENTH vehicle. Experimental results demonstrate the effectiveness of RSTP and highlight the impact of model composition weights on performance. Future work will extend RSTP to racing overtaking scenarios and incorporate constraints-in-the-loop.
@@
This appendix is structured as follows: Section -A details the experimental setup and hardware configurations; Section -B examines the kinematic feasibility of trajectories generated by the RSTP method; and Section -C evaluates the performance of the model composition within simulation.
-A Experiment
@@
To evaluate the RSTP, training and evaluation datasets are constructed from a real-world $6\,\text{m}\times 6\,\text{m}$ map. From the map’s free navigable area, a comprehensive evaluation dataset $D_{A}$ containing $N_{D}^{\text{all}}=1045$ uniformly sampled initial poses is created. To specifically assess the compositional model, a subset $D_{C}$ with $N_{D}^{\text{comp}}=79$ initial poses is selected from $D_{A}$. Target poses remain fixed. The trajectory horizon $L$ is set to 128 with 100 diffusion steps during training. Denoising employs the DDIM method with 8 steps, and the number of candidate trajectories $N_{\text{filter}}$ for the safety filter is set to 8. Training and inference are performed using an RTX 4060 GPU and a 4.9 GHz Intel(R) i7-12700 CPU with 48GB RAM.
-B On the Kinematic Feasibility of our RSTP method
@@
The proposed RSTP method also demonstrates excellent kinematic feasibility, as presented in Table I. Specifically, the M.TE of RSTP shows a substantial reduction of 47.23% compared to BLM. This significant improvement indicates that the diffusion model can effectively and implicitly learn vehicle kinematics from the dataset. In contrast, methods such as $\text{A}^{\star}$, $\text{RRT}^{\star}$, and $\text{FMT}^{\star}$, which neglect vehicle kinematic constraints, lead to considerably larger tracking errors. Optimization-based methods, such as RITP and OBCA, achieve notable reductions in M.TE of 70.67% and 79.27%, respectively, yielding trajectories with strong kinematic feasibility. However, this improved accuracy comes at a high computational cost. In comparison, RSTP generates trajectories with strong kinematic feasibility while demanding lower computational resources.
-C Effectiveness of Model Composition and Sensitivity Analysis of Compositional Weights
@@
The effectiveness of diffusion models composition for generalization is validated in scenes $CS_{1}$ (Fig. 6(a)) and $CS_{2}$ (Fig. 6(b)), employing four distinct compositional weight configurations, as summarized in Table III. As indicated in the table, the composed model can safely generalize to unseen scenes when appropriate compositional weights are selected, achieving an F.Rate of 0%. However, increasing the weight assigned to conditional models results in a rise in F.Rate across both test scenes. These results demonstrate that the generalization capability is sensitive to compositional weights, and maintaining appropriate weights is critical for minimizing failures.